# Bayesian Time-Varying Meta-Analysis via Hierarchical Mean-Variance Random-effects Models

### v1

@@ Introduction
Figure 2 shows box plots of the mean values for merchants and business types, calculated from the posterior distributions of the estimated merchant and business type effects. As shown in Figure 2, since the dataset comprises different experiments, there is between-study heterogeneity, and the effects vary depending on the merchant and business type. This result suggests that modeling random effects using a hierarchical Bayesian approach is appropriate. The coefficients for the business type effect are $-0.108$ for restaurants and $0.284$ for retail, with the value being higher for retail. These results indicate that the promotions encourage consumer stockpiling, as reported in previous studies (e.g. Ailawadi et al., 2007; Mela et al., 1998), and this effect may have appeared more prominently in the retail, where stockpiling is relatively easier. Figure 3 also shows the estimated time effects arranged chronologically. As shown in Figure 3, there is temporal heterogeneity in the dataset: the effect of promotions varies depending on the implementation period, with the highest effect in July and the lowest in April. These results suggest that marketers are likely to implement more effective promotions if they carry them out during the summer.

### v2

@@ Introduction
Figure 2 shows box plots of the mean values for merchants and business types, calculated from the posterior distributions of the estimated merchant and business type effects. As shown in Figure 2, since the dataset comprises different experiments, there is between-study heterogeneity, and the effects vary depending on the merchant and business type. This result suggests that modeling random effects using a hierarchical Bayesian approach is appropriate. The coefficients for the business type effect are 0.108-0.108 for restaurants and $0.284$ for retail, with the value being higher for retail. These results indicate that the promotions encourage consumer stockpiling, as reported in previous studies (e.g. Ailawadi et al., 2007; Mela et al., 1998), and this effect may have appeared more prominently in the retail, where stockpiling is relatively easier. Figure 3 also shows the estimated time effects arranged chronologically. As shown in Figure 3, there is temporal heterogeneity in the dataset: the effect of promotions varies depending on the implementation period, with the highest effect in July and the lowest in April. These results suggest that marketers are likely to implement more effective promotions if they carry them out during the summer.
@@
Competing interests
All authors report that NTT DOCOMO, INC provided financial support. Kohsuke Kubota and Keiichi Ochiai report a relationship with NTT DOCOMO, INC. that includes: employment. Shonosuke Sugasawa and Takahiro Hoshino report a relationship with NTT DOCOMO, INC. that includes: consulting or advisory.